OpenStudy (anonymous):

Find the slope of the line passing through the given points. (3,4) (2,-5/4) Tell wether the line rises, falls, is horizontal, or is vertical.

OpenStudy (precal):

use the slope formula

OpenStudy (radar):

If the slope is positive, the line rises. If the slope is negative the lines falls. Now the big question do you know what need to be done to calculate the slope with the information you have been provided by this problem?

OpenStudy (precal):

you can also, use graph paper and plot your 2 points to see the behavior of your line

OpenStudy (radar):

That too.

OpenStudy (radar):

If the line is horizontal the slope is zero (no slope) if the line is vertical the slope is undefined.

OpenStudy (radar):

Again what do you want to do?

OpenStudy (radar):

The graphing idea is the simplest and would provide visual re-enforcement!!.

OpenStudy (anonymous):

umm i know how to use the slope formula but i dont know how to do it in this case because one of the points is in fraction form.

OpenStudy (radar):

Let us give it a try the slope formula is (y2-y1)/x2-x1). Let us first label the pointsP1 is (3,4) and P2 is (2,-5/4) Now don't let that fraction bother you it is just a value. slope is m m=(-5/4-4)/(2-3) m=(-21/4)/-1 m=21/4 or 5 1/4 or 5.25 note it is positive. Your line will rise.

OpenStudy (anonymous):

okay.. i think i understand.. thanks

OpenStudy (radar):

you are welcome, graphing would of worked also.

OpenStudy (radar):

I just graphed it, yep it rises and rather steeply too. good day.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i just graphed it too.